The sum of first q terms of an A.P 63q-3q^2 if its pth term is -60 find the value of p also find the 5th term starting from backward pth term of this A.P

Grade:10

1 Answers

Arun
5 years ago
 
Dear student
 
T1 = S1 = 63-3 = 60
 
S2 = T1 + T2 = 126 – 12 = 114
hence T2 = 114- 60 = 54
 
Now you can solve . I hope
 
Regards
Arun